OpenStudy (fairy733189):

One solution of the equation 4x^2 - 9x + 5 = 0 is 5/4 . What is the other solution?

OpenStudy (fairy733189):

@Mehek14

jhonyy9 (jhonyy9):

so than 5/4 is a solution of this quadratic tis mean that this quadratic can being divided exactly witout remainder with (x-5/4) do you can dividing this quadratic by (x-5/4) ?

jhonyy9 (jhonyy9):

4x^2 -9x +5 = 0 one solution is 5/4 4x^2 -9x +5 / 4x-5 = x -1 4x^2 -5x --------- 0 -4x +5 -4x +5 -------- 0 0 so from this result that the second solution will be x-1=0 x=1 bc. this quadratic can be factorized 4x^2 -9x +5 = (4x-5)(x-1) hope this will nhelp you sure easy understanding the way how you solve same exercises good luck bye bye